Assume that the following cost data are for a purely competitive producer: 

Average Fixed Average Average Total Variable Cost Marginal Cost Total Product Cost Cost 000 S 0.00 na na 45.00 S 105.00 72.50 $ 45.00 60,00 S 42.50 S 40.00 30.00 S 6000 S 35.00 2000 S 40.00 S 52.50 S 30.00 15.00 $ 3760 S 3500 49.00 $ 12.00 S 37.00 S 1000$ 37.50 S 47.50 S 47.14 S 40.00 3857 S 45.00 857 $ 48.13 S 55.00 4063 S 7.50 $ 5000 $ 65.00 667 S 6.00 S 43 33 S 4650 S 52.50 S 75.00 10


Answer the questions in the first column in the table below for the price listed at the top of each of the other three columns 

Instructions: If you are entering any negative numbers be sure to include a negative sign () in front of those numbers. Select "Not applicable" and enter a value of "O for output if the firm does not produce At a product price of $68.00 At a product price of $43.00 At a product price of $34.00 Will this firm produce in the short run? Click to select) able to produce what vol be Theo can you etectunOA (Click to select)v output (Click to select -units output I-」units per firm output units profit-maximizing or loss-minimizing output? perf per firm What economic profit or loss will the firm realize per unit of output? Click to select per unit- S Click to select) per unit S 


d. In the table below, complete the shot.run supply schedule for the firm (columns 1 and 2) and indicate the profit or loss incurred at each output (column 3)

complete the short-run supply schedule for the firm (columns 1 and 2) and indicate the profit or loss incurred at each output (column 3) Instructions: Enter your answers as whole numbers. If you are entering any negative numbers be sure to include a negative sign Θ in front of those numbers. Quantity Supplied Single FirmProfit (+) or Loss (-) Quantity Supplied Price Firms S24.00 29.00 34.00 41.00 46 00 57.00 68 00


e. Now assume that there are 1,500 identical firms in this competitive industry, that is, there are 1,500 films, each of which has the cost data shown in the table. Complete the industry supply schedule (column 4 In the table above) 


f. Suppose the market demand data for the product are as follows: Price $24.00 17000 29 00 15000 34.00 13500 41 00 12000 46 00 10500 51.00 9500 56 00 8000 What will be the equilibrium price?$ What will be the equilibrium output for the industry units For each firm?

Instructions: Round your answers to 2 decimal places. Enter positive values for profit or loss. What will profit or loss be per unit? Click to select) per unit $ Per firm? $ Will this industry expand or contract in the long run? (Click to select)
Answer #1

a)

At a price of $68,

A perfectly competitive firm increases its output level as long as P>MC or P=MC.

We can see that MC is less than price for Q=9 but MC>P for Q=10

So, optimal output is 9 units.

At this output level AVC

As discussed above, profit maximizing or loss minimizing output is 9 units per firm.

Economic profit=(P-ATC)*Q=(68-50)*9=$162

b)

At a price of $43,

A perfectly competitive firm increases its output level as long as P>MC or P=MC.

We can see that MC is less than price for Q=6 but MC>P for Q=7

So, optimal output is 6 units.

At this output level AVC

As discussed above, profit maximizing or loss minimizing output is 6 units per firm.

Economic profit=(P-ATC)*Q=(43-47.50)*6=-$27 (Its a loss)

c)

At a price of $34,

We can see that AVC>P for each output level. So, firm will shut down in short run.

Hence, output of a firm=0

Economic Profit=Total Revenue at nil output-Total Cost at nil output=0-60=-$60 (Its a loss)

(Loss is equal to fixed cost in case of shut down)

d-e)

We can see that AVC>34 for every output level. So, firm will shut down for price less than 34 in the given case. Loss will be equal to fixed cost.

As described in part a and part b, firm will select its optimal output for Price higher prices. We can develop following schedule.

Price Q=Quantity supplied by single firm Profit (+)/Loss (-) Qs=Quantity suplied by 1500 firms=1500*Q
24 0 -60.00 0
29 0 -60.00 0
34 0 -60.00 0
41 6 (41-47.50)*8=-39.00 9000
46 7 (46-47.14)*7=-7.98 10500
57 8 (57-48.13)*8=70.96 12000
68 9 (68-50)*9=162.00 13500

f)

We can see that at price=$46, quantity demanded is equal to quantity supplied i.e. 10500. So,

Equilibrium price=$46

Equilibrium output for the industry=10500 units

For each firm=7 units

Profit per unit=(P-ATC)=46-47.14=-$1.14

(Its a loss of $1.14 per unit)

Per firm=(P-ATC)*Q=(46-47.14)*7=-$7.98

(Its a loss of $7.98 per firm)

Firm is making a loss in short run. Firm will start exiting the industry. So,

Industry will contract in the long run. (Contract)